OpenStudy (anonymous):

write an equation in slope -intercept form (if possible)for each line with a slope of -1/2 and y intercept 5

ganeshie8 (ganeshie8):

this is the slope intercept form : y = \(\color{red}{m}\)x + \(\color{green}{c}\)

ganeshie8 (ganeshie8):

where, slope = \(\color{red}{m}\) y intercept = \(\color{green}{c}\)

OpenStudy (anonymous):

so m = -1/2

ganeshie8 (ganeshie8):

right, and c = 5 just plugin :)

OpenStudy (anonymous):

so how do i set it up

ganeshie8 (ganeshie8):

y = mx + c plugin the given 'm' and 'c' values in above equation

ganeshie8 (ganeshie8):

y = -1/2x + 5

ganeshie8 (ganeshie8):

thats it !

OpenStudy (anonymous):

y=-1/2x+5

ganeshie8 (ganeshie8):

if u wondering why its called slope-intercept form, in slope-intercept form, we *see* slope and intercept directly

ganeshie8 (ganeshie8):

like this, there are few other forms for a line, which are useful in some particular situations

OpenStudy (anonymous):

so that would be my answer

ganeshie8 (ganeshie8):

ur previous reply is ur answer

OpenStudy (anonymous):

thanks

ganeshie8 (ganeshie8):

np :)